# How Far Can Transformers Reason? The Globality Barrier and Inductive Scratchpad

### v2

@@ Appendix E Comment on Lemma 1
Appendix F Discussion on circuit complexity connections
One approach we can use to analyse what we can learn with different methods is to consider the complexity class of the problems that can be solved by algorithms of a given type. Constant depth neural nets with well-behaved activation functions and weights of size at most polynomial in the input length are limited to computing functions in (possibly nonuniform) $TC^{0}$, the class of functions computable by polynomial-sized constant depth circuits built from AND, OR, NOT, and threshold gates. Likewise, constant depth Transformers with polynomial-sized weights, polynomial-sized alphabets, and attention matrices thats entries are rational with polynomial-sized numerators and denominators are limited to computing functions in $TC^{0}$.
@@ Appendix F Discussion on circuit complexity connections
Appendix G More experiments with ChatGPT

### v3

@@ Appendix G Comment on Lemma 1
Appendix H Discussion on circuit complexity connections
One approach we can use to analyze what we can learn with different methods is to consider the complexity class of the problems that can be solved by algorithms of a given type. Constant depth neural nets with well-behaved activation functions and weights of size at most polynomial in the input length are limited to computing functions in (possibly nonuniform) $TC^{0}$, the class of functions computable by polynomial-sized constant depth circuits built from AND, OR, NOT, and threshold gates. Likewise, constant depth Transformers with polynomial-sized weights, polynomial-sized alphabets, and attention matrices whose entries are rational with polynomial-sized numerators and denominators are limited to computing functions in $TC^{0}$.
@@ Appendix H Discussion on circuit complexity connections
Appendix I Experiments with ChatGPT